In a uniformly random permutation of the vowels {a,e,i,o,u}, what is the probability that the first letter is i and the last letter is o?
Answer #1

total number of arrangements of 5 vowels =5! =120

number of arrangements where first letter is i and the last letter is o =3! =6 (since only need to arrange middle 3 letters)

therefore probabiltiy =6/120 =1/20